OpenStudy (anonymous):

Let y = 5 sqrt(x). Find the change in y, ∆y when x= 4 and ∆x = 0.3

OpenStudy (anonymous):

∆y = f(x+∆x) - f(x) ∆y = 5( x + ∆x)^(1/2) - 5x^(1/2) This is as far as I got. Help please.

OpenStudy (freckles):

have you plugged in your given x and delta x?

OpenStudy (freckles):

the rest is plug in in and order of operations

OpenStudy (freckles):

\[\Delta y=5(4+0.3)^\frac{1}{2}-5(4)^\frac{1}{2}\]

OpenStudy (freckles):

@cjroyg are you still there?

OpenStudy (anonymous):

Sorry my computer totally crashed.

OpenStudy (anonymous):

But thank you for your assistance.

OpenStudy (freckles):

np